OpenStudy (anonymous):

Factor -2bk2 + 6bk - 2b.

mathslover (mathslover):

here a = -2b and b = 6b and c =-2b

mathslover (mathslover):

solve through quadratic equation : \[\huge{x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}}\]

OpenStudy (anonymous):

could you give the answer please

mathslover (mathslover):

so u r giving test right ?

OpenStudy (anonymous):

it is a quiz but i realy need to pass it

OpenStudy (anonymous):

-2bk2 +6bk-2b -2b(k^2 -3k +1) apply the quadratic formula the factors will be \[3\div2 +\sqrt{5}\div2\] \[3\div2 -\sqrt{5}\div2\]